OpenStudy (phebe):

MEDALSSSS!!!! 1. What is the factored form of q^2-12q+36 A-(q+6)(q-6) B-(q-6)(q-6) C-(q-9)(q+4) D-(q+4)(q+9)

OpenStudy (mathstudent55):

Check if it follows this pattern: \(a^2 - 2ab + b^2 = (a - b)^2\)

OpenStudy (anonymous):

B

OpenStudy (mathstudent55):

Also, \((a - b)^2 = (a - b)(a - b) \)

OpenStudy (phebe):

@mathstudent55 no it doesn't

OpenStudy (phebe):

@Ezack thanx

OpenStudy (mathstudent55):

Yes, it does. In your case, a = q and b = 6.

OpenStudy (anonymous):

(q-6)(q-6)=0 q^2-6q-6q+36=0 q^2-12q+36 = 0

OpenStudy (anonymous):

@Phebe no problem